# Evolving music theory for emerging musical languages

### v1

@@ Introduction
A foundational assumption in both psychoacoustics and Music Information Retrieval (M.I.R.) holds that harmonic complex tones in musical contexts convey a single pitch, typically corresponding to their fundamental frequency ($f_{0}$). This view underpins much of the theoretical discourse on music. Yet its universality is contested – both by historical debate, such as the Ohm-Seebeck controversy (Turner, , 1977), and by contemporary practice, including the use of multiphonics in acoustic (Fallowfield, , 2020) and electronic contexts (Deruty et al., 2025a, ; Deruty et al., 2025b, ; Deruty et al., 2025c, ; Deruty et al., 2025d, ). A single quasi-harmonic signal may give rise to multiple perceived pitches, with listeners potentially disagreeing on which are heard (Deruty et al., 2025c, ). This final observation, in particular, raises a central question: is pitch an ontologically objective property of a tone, or a listener-dependent construct?
@@
1 The framework of Western classical music
@@ 3.1 Specification 5: one harmonic tone, one pitch
Specification 5 has significant implications. Because harmonic tones are typically associated with a single perceived pitch, this pitch has often been treated – in Mari et al., ’s (2023) terms – as an ontologically objective property, meaning it is assumed to exist independently of any observer. This view reflects a naive realist approach to measurement, where the perceived pitch is taken as the ‘true’ value simply waiting to be uncovered. It forms the basis for most monophonic pitch-tracking algorithms (Drugman et al., , 2018; Kim et al., , 2018; Riou et al., , 2023). However, if perceived pitch can vary across listeners and listening conditions – as demonstrated by Deruty et al., 2025c in the case of synthetic quasi-harmonic tones – then the assumption of objectivity no longer holds. In the terms of Poincaré, (1902, p. 2), such a principle must instead be understood as a conventional hypothesis – that is, a methodological convenience rather than a universal truth.
@@ 3.2 Specification 9: two harmonic tones, two pitches
DeWitt and Crowder, (1987, p. 84) revisited Stumpf’s original findings and conducted their own experiments. While the authors consider their results strong enough to ‘elevate Stumpf’s principle of fusion from conjecture to established fact’, and therefore making Specification 9 a conventional hypothesis, the outcomes are not entirely consistent: quantitative measures of fusion vary across experiments. These discrepancies may stem from differences in experimental protocols (e.g., percentage of ‘one tone’ responses, mean reaction times) and in the acoustic properties of the stimuli (unspecified for Stumpf; triangle and square waves in later studies).
@@ 3.4 Tonal fission
DeWitt and Crowder, (1987, p. 84) conclude their study by suggesting that pitch perception ambiguities associated with tonal fusion could result from a single musical tone containing harmonics. This conjecture is supported by Deruty et al., 2025a ; Deruty et al., 2025b ; Deruty et al., 2025d ; Deruty et al., 2025c , who describe the deliberate use of synthetic quasi-harmonic tones to convey multiple and ambiguous pitches. A similar phenomenon has previously been observed in multiphonics produced from acoustic sources (Fallowfield, , 2020; Fox and Fallowfield, , 2020).

### v2

@@ Introduction
A foundational assumption in both psychoacoustics and Music Information Retrieval (M.I.R.) holds that harmonic complex tones in musical contexts convey a single pitch, typically corresponding to their fundamental frequency ($f_{0}$). This view underpins much of the theoretical discourse on music. Yet its universality is contested – both by historical debate, such as the Ohm-Seebeck controversy (Turner, , 1977), and by contemporary practice, including the use of multiphonics in acoustic (Fallowfield, , 2019) and electronic contexts (Deruty et al., 2025a, ; Deruty et al., 2025b, ; Deruty et al., 2025c, ; Deruty et al., 2025d, ). A single quasi-harmonic signal may give rise to multiple perceived pitches, with listeners potentially disagreeing on which are heard (Deruty et al., 2025c, ). This final observation, in particular, raises a central question: is pitch an ontologically objective property of a tone, or a listener-dependent construct?
@@
1 The Western classical music framework
for pitch transmission
@@ 3.1 Specification 5: one harmonic tone, one pitch
Specification 5 has significant implications. Because harmonic tones are typically associated with a single perceived pitch, this pitch has often been treated – in Mari et al., ’s (2023) terms – as an ontologically objective property, meaning it is assumed to exist independently of any observer. This view reflects a naive realist approach to measurement, where the perceived pitch is taken as the ‘true’ value simply waiting to be uncovered. It forms the basis for most monophonic pitch-tracking algorithms (Drugman et al., , 2018; Kim et al., , 2018; Riou et al., , 2023). However, if perceived pitch can vary across listeners and listening conditions – as demonstrated by Deruty et al., 2025c in the case of synthetic quasi-harmonic tones – then the assumption of objectivity no longer holds. In the terms of Poincaré, (1902, p. 2), such a principle must instead be understood as a disguised convention – that is, a methodological convenience rather than a universal truth.
@@ 3.2 Specification 9: two harmonic tones, two pitches
DeWitt and Crowder, (1987, p. 84) revisited Stumpf’s original findings and conducted their own experiments. While the authors consider their results strong enough to ‘elevate Stumpf’s principle of fusion from conjecture to established fact’, and therefore making Specification 9 a disguised convention, the outcomes are not entirely consistent: quantitative measures of fusion vary across experiments. These discrepancies may stem from differences in experimental protocols (e.g., percentage of ‘one tone’ responses, mean reaction times) and in the acoustic properties of the stimuli (unspecified for Stumpf; triangle and square waves in later studies).
@@ 3.4 Tonal fission
DeWitt and Crowder, (1987, p. 84) conclude their study by suggesting that pitch perception ambiguities associated with tonal fusion could result from a single musical tone containing harmonics. This conjecture is supported by Deruty et al., 2025a ; Deruty et al., 2025b ; Deruty et al., 2025d ; Deruty et al., 2025c , who describe the deliberate use of synthetic quasi-harmonic tones to convey multiple and ambiguous pitches. A similar phenomenon has previously been observed in multiphonics produced from acoustic sources (Fallowfield, , 2019; Fox and Fallowfield, , 2020).